OpenStudy (anonymous):

The slope-intercept form of a linear equation is y = mx + b, where x and y are coordinates of an ordered pair, m is the slope of the line, and b is where the line crosses the y-axis. Which is an equivalent equation solved for the slope, m?

OpenStudy (anonymous):

y2 minus y1 over x2 minus x1

OpenStudy (anonymous):

Thats the equation to find slopee

OpenStudy (anonymous):

m = yx + b m= y-b/x m= y/x-b m= y-b/x

OpenStudy (anonymous):

A, im almost positive

OpenStudy (shinalcantara):

y = mx + b y - b = mx m = (y-b)/x that is if we just have to utilize the y-intercept form

OpenStudy (anonymous):

oh shoot, yea @shinalcantara is correct